Question 1 Calculate the Coefficient of variation (percent) for a set of data where the mean (X) = 89 mg/dL and 2SD (s) = 14 mg/dL.

a. 7%

b. 7.8%

c. 7.9%

d. 14%

Question 2 A batch of test results is out-of-control. What should you do first?

a. Report the results to the physician first, and then look for the trouble.

b. Repeat the tests with a new lot of standards (calibrators).

c. Follow the "out of control" procedure specified for the test method.

d. Repeat the tests with a new lot of reagents.

Question 3 A new standard (calibrator) has been prepared in error at a lower concentration than that required for the test. How would such an error appear on a quality control chart?

a. Down ward shift

b. Upward shift

c. Down ward trend

d. Upward trend

Question 4 On a QC chart, when would a statistical out of control situation requiring corrective action be suspected?

a. When 6 successive plots fall above and below the mean within +/- 1s

b. When 6 successive plots fall above and below the mean within +/- 2s

c. When one plot falls within the area of +/- 2s to 3s within a 20 consecutive day span

d. When one plot falls outside the area of +/-3s within a 20 consecutive day span

Question 5 Which term refers to the measure of scatter of experimental data around the mean of a Gaussian (normal) distribution curve?

a. Mean

b. Coefficient of variation

c. Mode

d. Standard deviation
Answer #1

1. Coefficient variation in percent = SD/ mean x 100

Here 2SD = 14

SD = 14/2 = 7

Mean = 89

Coefficient variation = 7/89 x 100 = 7.86% or 7.9%

So the answer is (b) 7.9%

2. If the batch of test results is out of control then Follow the "out-of-control" procedure
specified for the test method.

So the answer is (c)

3. such an error appear on a quality control chart as upward shift.

So the answer is ( b)

4. The answer is (d)

5. Standard deviation refers to the measure of scatter of experimental data around the mean of a Gaussian distribution curve.